A gift grab bag is given to party goers. They are invited to reach into the grab bag and take one gift out. It contains 12 chocolates, 15 hard candies, 10 boxes of crackers, and 5 bags of cheese. Let H = the event of getting a chocolate. Let N = the event of getting a hard candy. Let F = the event of getting crackers. Let C = the event of getting a bag of cheese. What is P(N)?
Answer #1

Total number pf gifts = 12 + 15 + 10 + 5 = 42

Let N = the event of getting a hard candy.

Number of hard candies = 15

We want to find, P(N)

P(N) = 15/42 = 0.357143

=> P(N) = 0.3571 ( rounded to 4 decimal places)